If A(2,-3) and B(3,5) are the two vertices of a rectangle ABCD, find the slope of 
​1) BC 
​2) CD
​3) DA

Grade:12

1 Answers

Arun
6 years ago
Slope of AB = 5-(-3)/ 3-2 = 8/1 = 8
As we know for a reactangle side AB and CD will be
parallel
& Side BC & DA will be parallel.
Hence slope of CD= slope of AB = 8
 
Now BC will be perpendicular to AB.
Hence slope of BC = -1/8
So 
Slope of BC = slope of DA = -1/8